OpenStudy (kingj):

What is the sum of the arithmetic sequence 3, 9, 15…, if there are 34 terms?

OpenStudy (kingj):

so what is the answer

OpenStudy (kingj):

ok

OpenStudy (kingj):

these are the only aswers you can choose from 3,072 3,468 4,242 4,486

OpenStudy (kingj):

so im guessing number two is the best bet

OpenStudy (sasogeek):

d=6 a=3 \[\ S _{n} = \frac{n}{2}[2a+(n-1)d] \]\[\ S_{n}=\frac{34}{2}[2(3)+(34-1) \times 6] \]\[\ S_{n}=17[6+33 \times 6] \]

OpenStudy (anonymous):

^^sasogeek has the better explanation i guess I got too caught up in calculus ro figue out a simpler way

OpenStudy (anonymous):

*to figure (darn keyboard)

OpenStudy (anonymous):

I'm deleting my past answers to avoid confusion.

OpenStudy (sasogeek):

\[\ S_n=3468 \]